OpenStudy (anonymous):

a man travels a distance of 1.5 meter towards the east, then 2.0 meter towards the south and finally 4.5m towards the east in 4 s. Calculate (2)distance travelled and (b) displacement of the body Plz give full solution as i am in 9th class Need full solution plz

OpenStudy (sriram):

the total distance travelled by the body=1.5+2+4.5=8

OpenStudy (sriram):

the total displacement= length of line joining starting point and ending point in your Question if u imagine the path the body has finally travelled 1.5+4.5meters east =6 meters east & 2 meters south distance btwn starting starting and ending point =sqrt( 6^2 + 2^2 ) (pythagoras theorem) =sqrt(36+4) = sqrt(40)

OpenStudy (a_clan):

See the attachment